### v2

# Improving Sensitivity of Vector-like Top Partner Searches with Jet Substructure

Abstract
Vector-like quark partners appear in many BSM models and remain an important area of research, as they can offer insights into the electroweak symmetry breaking mechanism. In this work, we have focused on studying the production of a heavy vector-like top partner in association with a SM top quark via chromomagnetic coupling and the four decay modes of the top partner, namely, bW, tZ, th, and tg. The signal has been studied in final states with one fat jet, at least one b-tagged jet, one lepton, and missing energy. This study focuses on the extensive use of jet substructure techniques in jets clustered with fixed and dynamically varying radius to deal with events containing differently sized jets. Important kinematic information, along with jet substructure and event shape observables, has been used in a multivariate analysis to extract signal with high significance. A comparative study between fixed and dynamically varying radius clustering of jets is also presented, leading to an improvement in signal sensitivity in the highly boosted scenario.
preprint: IMSc/2025/04
I Introduction
The search for new physics beyond the standard model (BSM) remains one of the most important goals of current and future high-energy collider experiments. Despite many successes, the standard model (SM) of elementary particles leaves several fundamental questions unanswered, such as the nature of electroweak symmetry breaking and the stability of the electroweak vacuum Giudice:2013yca ; Arkani-Hamed:2012dcq ; Alekhin:2012py ; Degrassi:2012ry ; Dawson:2010jx , the origin of mass hierarchies Froggatt:1978nt ; Fritzsch:1999ee , matter-antimatter asymmetry Riotto:1999yt ; Dine:2003ax , the composition of dark matter Clowe:2006eq ; Sofue:2000jx , etc. Many BSM models Perelstein:2003wd ; Matsedonskyi:2012ym ; Chanowitz:1978mv propose the existence of new resonances, such as vector-like fermions Aguilar-Saavedra:2013qpa ; Falkowski:2013jya ; Dawson:2012di ; delAguila:1989rq , that can provide critical insights into some of these open problems and offer experimentally verifiable predictions at the TeV scale CMS:2024bni . Vector-like fermions are directly related to the stability of the electroweak vacuum Hiller:2022rla . A vector-like top partner, being a colour triplet and having the same electric charge as the SM top quark, couples to the Higgs boson and alters the renormalization group evolution of the Higgs quartic coupling. Its Yukawa interaction contributes positively to the running of the Higgs quartic coupling, while its mixing with the SM top quark reduces the Yukawa contribution due to the top quark that typically drives the coupling negative. Together, these effects slow down or prevent the Higgs quartic coupling from turning negative at high energy scales, thus stabilizing the electroweak vacuum all the way up to the Planck scale Hiller:2022rla . In addition, coloured vector-like fermions can mediate interactions between dark matter and SM quarks as indicated in the phenomenological analyses of vector-like portal dark matter models Ghosh:2024boo ; Colucci:2018vxz ; Ghosh:2025agw ; Ghosh:2022rta ; Ghosh:2023xhs . Thus, identifying such BSM states would not only signify the discovery of new physics phenomena but also provide a profound understanding of the structure of fundamental interactions.
Extraction of rare BSM signals from overwhelming SM backgrounds requires improved analysis techniques and theoretical modelling. State-of-the-art jet substructure methods Larkoski:2017jix ; Marzani:2019hun ; Kogler:2018hem , clustering algorithms Salam:2010nqg ; Sapeta:2015gee , and machine learning techniques Duarte:2024lsg ; Radovic:2018dip have become indispensable tools for identifying boosted heavy resonances and complex event topologies that are characteristic of various BSM signatures at the Large Hadron Collider (LHC). At the same time, access to high-quality data and simulation samples, such as CMS Open Data CMS_Open_Data ; CERN-OPEN-2020-013 , allows for a more reliable background model and realistic performance estimates.
The nature of modern BSM searches requires exploration at increasingly higher energy regimes. At these high energies, fat jets frequently arise from the hadronic decays of boosted heavy particles such as the top quarks or the heavy bosons, after parton showering and hadronization. One of the primary aspects of this paper is to assess the efficacy of the recently proposed “Dynamic radius” (DR) jet clustering algorithm Mukhopadhyaya:2023rsb , which is advantageous in the higher energy regime where such fat jets are prominent. The substructure of the fat jets originating from massive boosted particle decays differs significantly from that of narrow QCD jets Sterman:1977wj owing to their decay history. The DR algorithm leverages this substructure – more precisely, combinations of energy correlation functions Larkoski:2013eya to adaptively set the jet radius based on local kinematics, thus producing variable-sized jets in an event. Some recent studies have made use of the DR clustering algorithm with zero pile up scenario Kar:2022hxn ; Mukhopadhyaya:2023akv ; Ghosh:2025gue . However, understanding the performance of this algorithm in realistic experimental conditions remains extremely crucial.
In this work, we investigate the production of a heavy vector-like top partner ($T$) in association with a SM top quark, mediated through chromomagnetic interactions Buchkremer:2013bha ; Choudhury:2021nib ; Balaji:2021lpr . The top partner decays through four main channels: $\rm tZ$, $\rm th$, $\rm tg$, and $\rm bW$; each leading to distinct final state signatures. [^1] In this article, we followed a convention to denote all SM particles in roman and BSM particles in italics. However, the associated fields of both SM and BSM particles are denoted in italics. Among these, direct searches at the LHC have been performed primarily in the $\rm tZ$ and $\rm th$ channels, leading to stringent lower bounds on the mass of the top partner around a few TeV ATLAS:2024xne ; CMS:2023agg ; CMS:2024qdd . Furthermore, indirect constraints on the BSM nature of the $\rm tg$ coupling have been derived using $\mathrm{t\bar{t}}$ events Hayreter:2015ryk ; CMS:2022voq . However, only a handful of direct searches in the $\rm bW$ decay mode have been performed at the LHC with full Run-II data CMS:2022fck ; CMS-PAS-B2G-23-009 due to its limited sensitivity. The final state for this study consists of exactly one lepton, exactly one fat jet, at least one narrow jet that is b-tagged, and missing transverse momentum ($\text{p}_{\text{T}}^{\text{miss}}$), capturing event topologies sensitive to the presence of vector-like top partners. The bW decay mode of the $T$ quark constitutes a significant contribution to this specific final state. The probe for the signal process is conducted using jets clustered using traditional fixed-radius as well as the more recent DR approaches. We employ powerful jet substructure tools, namely Soft Drop (SD) grooming Larkoski:2014wba , $N$-subjettiness Thaler:2010tr , and energy correlation functions (ECFs) Larkoski:2013eya to better understand the nature and the behaviour of the fat jet in the event. In addition, we also studied several event shape variables, such as Sphericity Bjorken:1969wi ; Donoghue:1979vi , Aplanarity ATLAS:2012tch , Fox-Wolfram moments (HHFox:1978vw ; Bernaciak:2012nh etc., along with lepton and jet kinematic information in a multivariate analysis (MVA), to effectively distinguish signal events from the background. The signal events are simulated with MadGraph5_aMC@NLO Alwall:2014hca ; Frederix:2018nkq together with Pythia8 Bierlich:2022pfr for hadronization and parton showering and Delphes deFavereau:2013fsa for detector response modelling. For SM backgrounds, high statistics Monte Carlo (MC) samples from CMS Open Data have been utilized for a reliable estimate. This study also includes pile up simulation with a realistic value of the mean pileup interactions, considering the Run-II LHC conditions.
Before proceeding further, it should be noted that the main goal of this work is to explore the usefulness of the DR jet clustering in highly boosted final states and to compare its performance with standard fixed-radius clustering. Impact of various uncertainty sources, such as pile up contamination removal, jet energy scale and resolution, hadronization model etc., on the DR algorithm require dedicated in-depth studies that is beyond the scope of the present work. This study is intended as a proof-of-concept, focusing on qualitative improvements achievable with DR jet clustering using a purely statistical approach. This is an initial step to demonstrate the applicability of the DR jet clustering, considering a well-motivated candidate BSM process involving a heavy top partner coupled strongly with the SM top quark via chromomagnetic coupling.
The article is organized as follows. Section II.1 introduces the signal model based on chromomagnetic coupling between the vector-like top partner and the SM top quark, while the description of signal and background simulation procedures is presented in Sections II.2 and II.3, respectively. The DR algorithm and relevant jet substructure techniques are presented in Section III together with a control study using $\mathrm{t\bar{t}}$ + jets events. Section IV discusses the event reconstruction and selection strategy. In Section V, we detail the multivariate analysis strategy, optimization methods, and present our main results. Finally, Section VI provides a summary of the work and an outlook towards the future prospects.
II Signal and Background processes
II.1 Signal Model
The vector-like $T$ quark is a triplet under $SU(3)$, and its electric charge is determined by its hypercharge and isospin structure. Additionally, it can couple with any generation of the SM quark. We assume the heavy top partner predominantly interacts with the third-generation SM quarks. The $T$ quark is a singlet under weak isospin and has the same electromagnetic charge as the top quark; hence, it mixes with the top quark. The general Lagrangian terms are given below Buchkremer:2013bha ; Choudhury:2021nib ; Balaji:2021lpr .
$\mathcal{L}_{TWb}$ $=-\dfrac{g_{w}~\sin\theta_{L}}{\sqrt{2}}~\overline{T}~\gamma^{\mu}P_{L}bW_{\mu}+h.c.,$ (1)
$\mathcal{L}_{TZt}$ $=\dfrac{g_{w}}{2c_{w}}~\overline{t}~\gamma^{\mu}\left(2\sin\theta_{L}\cos\theta_{L}P_{L}+\sin\theta_{R}\cos\theta_{R}P_{R}\right)~TZ_{\mu}+h.c.,$ (2)
$\mathcal{L}_{Tht}$ $=-s_{w}\dfrac{\sin\theta_{R}\cos\theta_{R}}{2M_{W}}~\overline{t}~\left(M_{T}P_{L}+m_{t}P_{R}\right)~T~h+h.c.$ (3)
Here, $t$ and $h$ denote the SM top quark and higgs boson, respectively. $P_{L/R}=\dfrac{1\mp\gamma_{5}}{2}$ represent the left- and right-chiral projection operators. The mixing angle between the left-chiral (right-chiral) SM top quark and the left-chiral (right-chiral) $T$ quark is denoted by $\theta_{L}$ ($\theta_{R}$). The symbols $M_{T}$, $m_{t}$, and $M_{W}$ represent the masses of the $T$ quark, the SM top quark, and the WW boson, respectively. The $SU(2)_{L}$ gauge coupling is $g_{w}$; $s_{w}$ and $c_{w}$ are sine and cosine of the electroweak mixing angle. The single production of the exotic $T$ quark in association with the SM top quark, is possible at the LHC via gluon fusion. The effective Lagrangian of such interactions is as follows.
$\mathcal{L}_{tg}=\dfrac{g_{s}}{\Lambda}~G^{a}_{\mu\nu}~\left[C_{tRL}\overline{T}\sigma^{\mu\nu}\tau^{a}P_{L}~t+C_{tLR}\overline{T}\sigma^{\mu\nu}\tau^{a}P_{R}~t\right]+h.c.$ (4)
where $\sigma_{\mu\nu}=i/2~(\gamma_{\mu}\gamma_{\nu}-\gamma_{\nu}\gamma_{\mu})$, and $g_{s}$ is the SM strong coupling constant. The notation $G^{a}_{\mu\nu}$ represents the gluon field strength tensor, and $\tau^{a}$ corresponds to the $SU(3)$ generators. The notation $\Lambda$ denotes the new physics scale, and $C_{tRL}$, $C_{tLR}$ are the coupling strengths. Corresponding to the above Lagrangian given in Eqs. (1)–(4), the $T$ quark has four decay channels, namely bW, tZ, th, and tg. While the Eq. (4) primarily contributes to the production of the $T$ quark, it is also responsible for the tg decay mode. From Eqs. (1)–(3), one notices that the branching ratios for $T\to\rm tZ$, $T\to\rm bW$, and $T\to\rm th$ become negligibly small when the mixing angles are very small ($\theta_{L},\theta_{R}\to 0$), and the $T$ quark decays predominantly into a top quark and a gluon. However, with increasing mixing between the $T$ quark and the top quark, the branching fractions into the electroweak channels $T\to\rm tZ$, $T\to\rm bW$, and $T\to\rm th$ increase significantly.
II.2 Signal Simulation
The process we are interested in is the associated production of an SM top quark and its vector-like partner $T$ via the effective chromomagnetic interaction.
$\mathrm{pp}$ $\to$ $\mathrm{t}\,\overline{T}+X,\quad{\rm and}$ (5)
$\mathrm{pp}$ $\to$ $T\,\overline{\rm t}+X.$ (6)
A set of representative parton-level Feynman diagrams corresponding to the process has been given in Figure 1.
Figure 1: Representative parton-level Feynman diagrams for the process $\mathrm{pp}\to T\,\bar{\mathrm{t}}\,/\,\bar{T}\,\mathrm{t}$.
To simulate the processes indicated by Eqs. (5) and (6), we first implemented all the interaction terms defined in Eqs. (1)–(4), along with the SM Lagrangian, using FeynRules Christensen:2008py ; Alloul:2013bka to generate a Universal FeynRules Output (UFO) Degrande:2011ua model file. This UFO model was then used in the MadGraph5_aMC@NLO framework to generate parton-level events at a centre-of-mass energy of $\sqrt{s}~=~13$~TeV with the parton distribution function (PDF) set to NNPDF30_nlo_nf_5_pdfas PDFset NNPDF:2014otw . The factorization and renormalization scales are set to the default dynamical scale in MadGraph5, which corresponds to the transverse mass of the $2\to 2$ scattering process at leading order (LO). Figure 2 shows the variation of the partonic cross section for $T$ quark production in association with a top quark at $\sqrt{s}~=~13$~TeV, along with the branching ratios of its various decay channels, for a fixed new physics scale of $\Lambda=8$ TeV.
Figure 2: Cross section vs. coupling (upper left) and vs. $T$ quark mass (upper right) are presented for $\mathrm{pp}\to T~\bar{\mathrm{t}}~/~\bar{T}~\rm t$ process at $\sqrt{s}~=$~13 TeV with new physics scale $\Lambda~=~8$~TeV, $M_{T}~=~2.2$~TeV (upper left), and $C_{tRL}~=~C_{tLR}~=~0.8$ (upper right). Branching ratios (BRs) vs. mixing angle for all $T$ decay modes (lower) are shown with $M_{T}~=~2.2$~TeV, $C_{tRL}~=~C_{tLR}~=~0.8$, and $\Lambda~=~8$~TeV.
To illustrate the collider analysis for this study, we consider a benchmark point that is consistent with existing experimental bounds. The strongest lower bound on $M_{T}$ is about 1.8 TeV ATLAS:2024xne . We further consider $\theta_{L}=\theta_{R}$ and $C_{tRL}=C_{tLR}$ to reduce the number of free parameters. The couplings $C_{tRL}$ and $C_{tLR}$ control the top-partner pair production cross section, which is restricted to $\mathcal{O}(1~{\rm fb})$ CMS:2024bni , resulting in an upper limit on $\frac{C_{tRL}}{\Lambda}=\frac{C_{tLR}}{\Lambda}\lesssim 0.6~\text{TeV}^{-1}$. Furthermore, the mixing parameters $\theta_{L}$ and $\theta_{R}$ affect the Higgs boson branching ratio to gluon-gluon and di-photon modes. One finds $\theta_{L}=\theta_{R}\lesssim 0.25$ radian for $M_{T}=2.2$ TeV from Refs. ATLAS:2025hhd ; Dawson:2012di ; Zhang:2025zkn . Based on these factors, we choose the benchmark scenario as follows:
$\frac{C_{tRL}}{\Lambda}=\frac{C_{tLR}}{\Lambda}=0.1~\text{TeV}^{-1},\qquad\theta_{L}=\theta_{R}=0.06~\text{radian},\qquad\text{and}\quad\ M_{T}=2.2~{\rm TeV}.$ (7)
The parton-level kinematics of the produced top partner are shown in Fig. 3 for different mass hypotheses. The higher tail of the $\text{p}_{\text{T}}$-spectrum of the $T$ quark has a distinct dependence on its mass. The cross section of the signal process, the $T$ quark decay width, and its branching fractions for all of the four decay modes are listed in Appendix A for different mass hypotheses.
Figure 3: Distributions of the top partner $\text{p}_{\text{T}}$ (left) and $\eta$ (right) at parton level for different mass hypotheses for the process $\mathrm{pp}\to T\,\bar{\mathrm{t}}\,/\,\bar{T}\,\mathrm{t}$ at $\sqrt{s}=13$ TeV.
The generated events are then passed through Pythia8 Bierlich:2022pfr for parton showering and hadronization. In the Pythia8 simulation, the CMS Tune MonashStar CMS:2015wcf , based on Monash 2013 tune Skands:2014pea , has been used to better match CMS data, especially to model the underlying event, initial-state radiation, and beam remnants. We then incorporated the effects of pile up in the signal MC samples by overlaying minimum-bias event samples with the hard-scattered events with approximately $20$ average pile up interactions per event. Finally, a fast detector simulation is performed on the pile up overlaid samples using Delphes_v3.5.0 deFavereau:2013fsa that employs an overall detector geometry definition and a parametric model of average detector response [^2] The choice of using Delphes fast detector simulation for signal MC samples was made owing to limitations in the available computing and storage resources at our disposal for this study.. The signal event simulation details are summarized in Table 1.
Table 1: Details of signal simulation tools and associated settings.
BSM model SM + Eqs. (1)–(4) using FeynRules Christensen:2008py ; Alloul:2013bka and
Universal FeynRules Output (UFO) Degrande:2011ua
Matrix element Eqs. (5) and (6) using MadGraph5_aMC@NLO Alwall:2014hca
at LO with $\sqrt{s}=13$ TeV
PDF set NNPDF30_nlo_nf_5_pdfas NNPDF:2014otw
Parton showering & Hadronization Pythia8 Bierlich:2022pfr
Underlying event CMS Tune MonashStar CMS:2015wcf , based on Monash
2013 tune Skands:2014pea
Detector simulation model Delphes_v3.5.0 deFavereau:2013fsa
Depending on the decay mode of the top partner into $\rm tZ$, $\rm th$, $\rm tg$, or $\rm bW$, the above processes could be realized into different event categories. To suppress a huge QCD multijet background at the LHC, we focus on the semi-leptonic channels requiring one lepton ($\ell\ =\ e,\mu$) in the final state. Therefore, the final states where one of the $\rm W$’s originating from either of the $\rm t$, $\rm{\bar{t}}$, $T$, or $\bar{T}$ decay leptonically, are studied. The event sample for our search can be broadly classified into the following two categories:
${\rm Two\ fat\ jets\ category:}$ $1\ell+2\ {\rm fat\ jets}+1\ {\rm b-jet}+\text{p}_{\text{T}}^{\text{miss}},$ (8)
${\rm One\ fat\ jet\ category:}$ $1\ell+1\ {\rm fat\ jet}+\geqslant 1\ {\rm b-jet(s)}+\text{p}_{\text{T}}^{\text{miss}}.$ (9)
Table 2 lists the different decay modes of the top partner and their possible contribution to each event category applying the same selection criteria as defined in Section IV on the lepton and jets. The signal fractions are larger in the two fat jets event category for all decay modes. However, in this event category, the $\mathrm{t\bar{t}}$ + X (X = W, Z, jets, etc.) backgrounds also have a significantly larger contribution, leading to a poor signal-to-background ratio and thus making it unsuitable as a search region for the signal process. Therefore, we focus on the one fat jet category where the contribution of the $\rm bW$ decay mode is relatively higher.
[…]
Decay mode
$T\to\mathrm{tZ}$ $T\to\mathrm{th}$ $T\to\mathrm{tg}$ $T\to\mathrm{bW}$
[…]
Two fat jets category $73.00\%$ $72.03\%$ $76.11\%$ $65.89\%$
One fat jet category $27.00\%$ $27.97\%$ $23.89\%$ $34.11\%$
II.3 Background Processes
[…]
accuracy [pb] accuracy [Million]
$\mathrm{t\bar{t}}$ + jets ttxj_full NLO Alwall:2014hca ; Frederix:2012ps ; Sjostrand:2007gs 833.900 Czakon:2011xx NNLO + NNLL 117.70
[…]
$\mathrm{t\bar{t}Z}$ ttZjets_full LO Alwall:2014hca ; MLM01 ; Mangano:2006rw ; Sjostrand:2007gs 0.863 Kulesza:2018tqz NLO + NNLL 22.16
$\mathrm{t\bar{t}W}$ ttWjets_full NLO Alwall:2014hca ; Frederix:2012ps ; Sjostrand:2007gs 0.745 Buonocore:2023ljm NNLO (QCD) 0.83
+ NLO (EW)
WZ + jets WZjets_full NLO Alwall:2014hca ; Frederix:2012ps ; Sjostrand:2007gs 56.850 Grazzini:2016swo NNLO 13.19
WW + jets WWjets_full NLO Nason:2013ydw ; Alioli:2010xd ; Sjostrand:2007gs 1.371 Grazzini:2016ctr NNLO 8.92
QCD multijet QCD_full LO Sjostrand:2007gs $1.981\times 10^{9}$ LO 97.19
Table 4: Stitching weights for $\mathrm{t\bar{t}}$ + jets events in different $\text{H}_{\text{T}}$ bins for 300 fb-1 integrated luminosity, computed following the prescription in Ref. Ehataht:2021rkh .
$\text{H}_{\text{T}}$ bins [GeV]
[0, 600) [600, 800) [800, 1200) [1200, 2500) $\geqslant 2500$
Stitching weights $2.562$ $2.550$ $1.413$ $9.180$ $2.154$
$\times 10^{-2}$ $\times 10^{-2}$ $\times 10^{-3}$ $\times 10^{-4}$
III Jet Algorithm and Substructure Methods
III.1 Dynamic Radius Jet Clustering Algorithm
The jet clustering algorithms are crucial steps in the analysis of events in a hadron collider experiment such as the LHC Salam:2010nqg ; Sapeta:2015gee . The currently used algorithms are fixed-radius sequential recombination inclusive jet algorithms, namely the $\text{k}_{\text{T}}$ (KT), anti-$\text{k}_{\text{T}}$ (AK), and Cambridge-Aachen (CA) Ellis:1993tq ; Catani:1993hr ; Cacciari:2008gp ; Dokshitzer:1997in ; Wobisch:1998wt algorithms. These traditional algorithms take as input the list of four momenta of the particles from an event and successively recombine them into several jets. In their iterative steps, these algorithms calculate pair-wise $\text{k}_{\text{T}}$” distance ($\text{d}_{\text{ij}}$) between $\text{i}^{\text{th}}$ and $\text{j}^{\text{th}}$ four momenta and “beam” distance ($\text{d}_{\text{iB}}$) for the $\text{i}^{\text{th}}$ four momentum, defined as
$\text{d}_{\text{ij}}=\min\left(\text{p}_{\text{T,\ i}}^{2n}\ ,\ \text{p}_{\text{T,\ j}}^{2n}\right)\ \frac{\Delta\text{R}_{\text{ij}}^{2}}{\text{R}^{2}},\ \text{and}$ (10)
$\text{d}_{\text{iB}}=\text{p}_{\text{T,\ i}}^{2n}$ (11)
to decide whether to merge two four momenta or declare a particular four momentum as a final jet. Here, $\Delta\text{R}_{\text{ij}}$ is the distance between $\text{i}^{\text{th}}$ and $\text{j}^{\text{th}}$ four momenta in the pseudorapidity-azimuth ($\eta$-$\phi$) plane. The two four momenta are merged to give a new four momentum if $\text{d}_{\text{ij}}$ becomes the least among all the pair-wise $\text{k}_{\text{T}}$” and beam distances. Otherwise, the $\text{i}^{\text{th}}$ four momentum corresponding to the least beam distance is declared as a final jet. Importantly, the algorithm has parameters $n$, which usually takes values 1, 0,and 1-1,\ 0,\ $\text${and}\ 1 corresponding to AK, CA, and KT algorithms respectively, and R as the distance measure in the $\eta$-$\phi$ plane. These algorithms typically output jets of a fixed size, which is determined by the input parameter R to the algorithm. In this article, the AK algorithm is chosen as the default since AK jets have regular conical shape with an well-defined jet area Cacciari:2008gp .
The traditional jet clustering approaches mentioned above have the limitation in producing varying-sized jets in an event for a given value of R. Several interesting attempts have been made to modify the algorithms to accommodate jets of variable sizes; such as “Variable R” jets Krohn:2009zg , “Fuzzy” jets Mackey:2015hwa , “SHAPER” Ba:2023hix , “XCone” Stewart:2015waa , “PAIReD” jets Mondal:2023law , “dynamic radius” jets Mukhopadhyaya:2023rsb etc. Here, we focus on the “dynamic radius” jet clustering algorithm Mukhopadhyaya:2023rsb , which attempts to dynamically modify the jet radius during the evolution of each jet through the iterative steps of the algorithm. The approach is to start from an initial radius $\rm R_{0}$}, the radius for each evolving “proto-jet” gets modified depending on its internal structure during the process of clustering. The square of the additive modifier $\sigma_{i}$ to $\rm R_{0}$} for each evolving “proto-jet” at each interactive stage is given by
$\sigma_{\rm i}^{2}=\dfrac{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}\ \Delta\mathrm{R}_{ab}^{2}}{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}}-\left(\dfrac{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}\ \Delta\mathrm{R}_{ab}}{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}}\right)^{2},$ (12)
where $a$ and $b$ run over the constituents of the proto-jet. So, the final radius for the $i^{\rm th}$ jet becomes $\mathrm{R_{d}}=\mathrm{R_{0}+\sigma_{i}}$. The advantage one gets in this method is that the radius is no longer fixed; it gets adapted to the requirement of the jet that is being formed. The modifier is a slight variation of a combination of a few jet substructure observable ECFs Larkoski:2013eya , which tends to become larger for top or heavy boson jets ATLAS:2019kwg . On the other hand, it tends to be smaller for narrow jets like the ones appearing from light quarks or gluons. Since our study involves one fat jet along with narrow b-jets, the efficiency of the DR jet clustering is expected to be better.
III.2 Jet Substructure and Parameter Choices
Improving signal sensitivity over the backgrounds using jets and their substructures is the primary aim of this work. We present here a brief overview of the various jet substructure methods that are used in our analysis.
Soft Drop Grooming
To mitigate contamination from pile up, underlying events etc., inside a jet, and improve the reliability of the jet substructure measurements, we have used the Soft Drop de-clustering method Larkoski:2014wba . This method primarily aims to remove soft and wide-angle radiations inside a jet while retaining the hard core of the jet. To briefly describe, the Soft Drop algorithm is applied on a jet that has been re-clustered using the CA algorithm (which orders by angle). The essential step of this method is to remove soft and wide-angle subjet after de-clustering the jet into two subjets (labelled as i and j) by checking the following condition.
$\mathrm{\frac{\min(p_{T,i},p_{T,j})}{p_{T,i}+p_{T,j}}}<z^{\rm SD}_{\text{cut}}\left(\mathrm{\frac{\Delta R_{ij}}{{\rm R}^{\rm SD}}}\right)^{\beta^{\rm SD}}$ (13)
Here, $\mathrm{p_{T,i}}$ and $\mathrm{p_{T,j}}$ are the transverse momenta of the subjets, $\mathrm{\Delta R_{ij}}$ is the distance between them in the $\eta$-$\phi$ plane. This process is repeated until the condition fails to be satisfied. The radius ${\rm R}^{\rm SD}$ is the characteristic jet radius usually taken to be the radius parameter used during the clustering. In case of the DR jet clustering algorithm, we have taken this parameter to be $\mathrm{R_{d}=R_{0}+\sigma}$, the final dynamically adjusted radius for each jet. The two parameters of Soft Drop are (1) $z^{\rm SD}_{\text{cut}}$: the energy fraction threshold and (2) $\beta^{\rm SD}$: the exponent regulating the angular separation between the two subjets. In the special case of $\beta^{\rm SD}=0$, the method reduces to the Modified Mass-Drop Tagger Butterworth:2008iy ; Dasgupta:2013ihk . The grooming then depends only on the momentum fraction $z$ and not on the angular separation. We have used the value of $\beta^{\rm SD}=0$ to keep our analysis aligned with the available jet substructure measurement ATLAS:2019kwg . We also check the alternative value of $\beta^{\rm SD}=1$. For the energy fraction threshold parameter, a nominal value of $z_{\rm cut}^{\rm SD}=0.1$ is used to match the studies performed in Refs. Larkoski:2014wba ; Larkoski:2014bia ; Larkoski:2015lea ; CMS:2017qlm .
NN-subjettiness
The jet substructure method $N$-subjettiness Thaler:2010tr , provides a measure of the consistency of a jet having $N$ subjets inside. In our event topology, we expect boosted top, W or Z fat jet, which gives rise to a three- or two-prong structure. Therefore, this method becomes particularly useful with the value of $N=3$ and 2. The $N$-subjettiness observable captures these subjets by fitting $N$ subjet axes inside the fat jet. This is performed by minimizing the $N$-subjettiness measure Thaler:2011gf defined as
$\mathrm{\tilde{\tau}_{N}^{(\beta^{\rm NS})}=\frac{1}{d_{0}}\sum_{i\in{\rm jet}}{p_{T}}_{i}}\min\left\{\Delta{\rm R}_{{\rm i},a_{1}}^{\beta^{\rm NS}},\Delta{\rm R}_{{\rm i},a_{2}}^{\beta^{\rm NS}},\cdots,\Delta{\rm R}_{{\rm i},a_{N}}^{\beta^{\rm NS}}\right\},$ (14)
where the sum runs over all constituents in the jet, and $\Delta{\rm R}_{{\rm i},a_{k}}$ is the distance in $\eta$-$\phi$ plane between constituent i and subjet axis $a_{k}$. The prefactor $\rm 1/d_{0}$} is a normalization factor. The minimization is performed over the choices of axes, and the final minimum value of $\tilde{\tau}_{N}$, denoted as $\tau_{N}$, represents the optimal measure of $N$-subjettiness. The exponent $\beta^{\rm NS}$ and the method used to determine the seed $N$ axes (e.g., exclusive $\text{k}_{\text{T}}$ axes from an iterative algorithm) are parameters of choice. We have used $N$-subjettiness method implemented in Fastjet Contrib Cacciari:2011ma . In our study, we use $\beta^{\rm NS}=1$, the seed axes obtained by WTA_KT_Axes method Bertolini:2013iqa ; Larkoski:2014uqa ; Salam_unpub , and the ratios $\tau_{32}=\tau_{3}/\tau_{2}$ and $\tau_{21}=\tau_{2}/\tau_{1}$ to respectively identify three- and two-prong structures inside a fat jet.
Energy Correlation Functions
These are a class of infrared and collinear safe observables proposed to study jet properties Larkoski:2013eya . The first three lowest point functions take the form
${\rm ECF}_{1}(\beta^{\rm ECF})$ $=$ $\mathrm{\sum_{i\,\in\,jet}p_{T,i}}\ ,$ (15)
${\rm ECF}_{2}(\beta^{\rm ECF})$ $=$ $\mathrm{\sum_{i<j\,\in\,jet}p_{T,i}\,p_{T,j}\,\Delta R_{ij}^{\beta^{\rm ECF}}}\ ,$ (16)
${\rm ECF}_{3}(\beta^{\rm ECF})$ $=$ $\mathrm{\sum_{i<j<k\,\in\,jet}p_{T,i}\,p_{T,j}\,p_{T,k}\left(\Delta R_{ij}\,\Delta R_{ik}\,\Delta R_{jk}\right)^{\beta^{\rm ECF}}}\ ,$ (17)
where the sum is over the jet constituents and the exponent $\beta^{\rm ECF}$ is a parameter that gives the weight to the angular separation between the jet constituents. These functions help distinguish between jets originating from different partonic configurations. In case of top quark study, a better discrimination from QCD jets is obtained for $\beta^{\rm ECF}=1.0$Larkoski:2013eya . We therefore choose to use $\beta^{\rm ECF}=1.0$ as a nominal value and $\beta^{\rm ECF}=0.5$ for an alternate value. Instead of using the ECFs as defined in Eqs. (15)–(17) directly, their ratios are better suited as they remove the $\text{p}_{\text{T}}$ dependence. The following ratios ATLAS:2019kwg are studied.
$\quad\ e_{2}=\displaystyle\frac{{\rm ECF}_{2}}{{(\rm ECF}_{1})^{2}}\ ,\qquad\qquad$ $e_{3}=\frac{{\rm ECF}_{3}}{({\rm ECF}_{1})^{3}}\ ,$ (18)
$\quad C_{2}=\displaystyle\frac{e_{3}}{(e_{2})^{2}}\ ,\qquad\text{and}\ \qquad\$$D_{2}=\frac{e_{3}}{(e_{2})^{3}}\ .$ (19)
The DR jet algorithm, along with the jet substructure methods, has been implemented within Delphes and CMS SoftWare version 7_6_7 as appropriate plugins. The choices of parameters for the jet substructure methods described above for the DR jets, including both nominal and alternate values used for variations, are summarized in Table 5. The impact of these choices is studied by comparing the shapes of different observables sensitive to the respective substructure methods, as shown in Figure 4, both at truth and reconstructed levels. As seen in the figure, the distributions for nominal and alternate parameter choices look quite similar, suggesting that the overall analysis remains fairly robust against these variations. In addition, substructure observables of the DR jets are also compared with the fixed-radius counterpart at the reconstructed level, and are shown in Figure 5. Only the nominal values of the substructure parameters are considered for this comparison. Figure 5 demonstrates that the distributions of $\tau_{21}$ and $\tau_{32}$ for the DR clustering are more shifted towards 0 compared to the fixed-radius clustering, indicating better reconstruction of the hadronically decaying boosted W bosons and top quarks in the former case with a fat jet Thaler:2010tr .
Table 5: Parameter choices for jet substructure methods applied to the DR fat jets. “Nominal Values” correspond to those used in the final analysis, while “Alternate Values” are considered for the control study shown in Figure 4.
Method Parameter Nominal Alternate
Values Values
Soft Drop Larkoski:2014wba $\beta^{\text{SD}}$ 0.0 ATLAS:2019kwg 1.0
$z_{\text{cut}}^{\text{SD}}$ 0.1 ATLAS:2019kwg 0.1
NN-subjettiness Thaler:2010tr $\beta^{\text{NS}}$ 1.0 ATLAS:2019kwg 1.0
  Seed Axes Stewart:2015waa   WTA_KT_Axes Larkoski:2014bia   WTA_KT_Axes
ECF Larkoski:2013eya $\beta^{\text{ECF}}$ 1.0 ATLAS:2019kwg 0.5
Figure 4: Comparisons of Soft Drop mass (upper-left), $\tau_{21}$ (upper-right), $\tau_{32}$ (middle-left), and the ECF ratios $C_{2}$ (middle-right), $D_{2}$ (lower-left), and $e_{3}$ (lower-right) distributions of the leading anti-$\text{k}_{\text{T}}$ fat jet corresponding to nominal and alternate values of substructure parameters as indicated in Table 5 for dynamic radius jet clustering algorithm at truth and reconstructed levels in $\mathrm{t\bar{t}}$ + jets events.
Figure 5: Comparisons of Soft Drop mass (upper-left), $\tau_{21}$ (upper-right), and $\tau_{32}$ (lower) distributions of the reconstructed leading anti-$\text{k}_{\text{T}}$ fat jet between fixed-radius (${\rm R}=0.8$) and dynamic radius jet clustering algorithms in $\mathrm{t\bar{t}}$ + jets events are presented. The nominal choice of substructure parameters as indicated in Table 5 has been used for the dynamic radius algorithm to obtain these plots.
IV Event Reconstruction and Selection
Various key features of the event reconstruction and selection are presented below.
Selection Criteria
Events having a minimum of one reconstructed pp interaction vertex are retained for this study. The vertices must be reconstructed from at least four tracks that have a longitudinal distance $|\text{d}_{\text{z}}|<24$ cm and a radial distance $|\text{d}_{\text{xy}}|<2$ cm from the nominal interaction point. If multiple vertices are found in an event, the one with the highest value of summed $\text{p}_{\text{T}}^{2}$ of clustered particles is taken as the primary pp interaction vertex. In contrast, the others are identified as the pile up vertices. Individual particles are reconstructed via the particle-flow algorithm CMS:2017yfk . Muons (Electrons) are required to have $\text{p}_{\text{T}}\geqslant 30\ (35)$ GeV, $|\eta|\leqslant 2.4\ (2.1)$ and to pass the “tight” identification criteria CMS:2018rym ; CMS:2020uim . The selected muon and electron candidates must pass the criterion on relative isolation ($\text{I}_{\text{rel}}$) to be within $6\%$ of their respective transverse momentum. Additionally, the selected electron candidate is required to satisfy $|\text{d}_{\text{xy}}|\leqslant 0.05\ (0.1)$ cm and $|\text{d}_{\text{z}}|\leqslant 0.1\ (0.2)$ cm relative to the primary vertex in the barrel (endcap) region of the electromagnetic calorimeter (ECAL). Events containing additional muons (electrons) with $\text{p}_{\text{T}}\geqslant 10\ (15)$ GeV, $|\eta|\leqslant 2.4\ (2.5)$, and passing the “loose” (“veto”) identification criteria are rejected. In such cases, the criteria on the additional lepton isolation are relaxed to $\text{I}_{\text{rel}}\leqslant 20\%$ for muons and $\text{I}_{\text{rel}}\leqslant 18\%\ (16\%)$ for electrons in the barrel (endcap) ECAL region. Standard jets are reconstructed using the AK clustering algorithm Cacciari:2008gp with radii of 0.4 and 0.8 for the narrow (AK4) and fat (AK8) jets respectively, as implemented in the FastJet package Cacciari:2011ma with the particle-flow candidates as inputs. The effect of additional tracks and energy deposits in calorimeters from pile up on the AK4 and AK8 jet momenta are mitigated by discarding tracks identified to be originating from pile up vertices, as well as by applying an offset correction to account for residual neutral pile up contributions Cacciari:2007fd ; CMS:2020ebo . “Tight lepton veto” identification criteria CMS-PAS-JME-16-003 are applied to suppress jets arising from spurious sources, such as electronic noise in the calorimeters and jets containing “tight”-ly identified muons or electrons. The combined secondary vertex tagging algorithm version 2 (CSVv2CMS:2017wtu is used to identify AK4 jets originating from bottom quarks or antiquarks. Candidate events are required to have exactly one AK8 jet with $\text{p}_{\text{T}}$ $\geqslant 200$~GeV and within $|\eta|\leqslant 4.0$, as mentioned in Section II.2. In addition to the selected AK8 jet, events must have at least one non-overlapping AK4 jet with $\text{p}_{\text{T}}\geqslant 30$ GeV, $|\eta|\leqslant 2.5$ and passing the “medium” working point defined on the CSVv2 score. Finally, events should pass $\text{p}_{\text{T}}^{\text{miss}}\geqslant 30$~GeV criterion.
We note that a full detector simulation with Geant4 with CMS detector specification has been used in the background MC samples. However, for the signal MC samples, a fast detector simulation using Delphes has been used. According to Ref. deFavereau:2013fsa , the overall agreement in the average performance of Delphes with the full simulation of the CMS detector based on Geant4 is reasonably good for different high-level physics objects, especially in the phase space region of interest. Furthermore, the measured efficiencies of lepton (both selected and vetoed leptons) CMS:2018rym ; CMS:2020uim and jet CMS-PAS-JME-16-003 ; CMS:2017wtu reconstruction are incorporated into the Delphes for signal events in bins of their respective $\text{p}_{\text{T}}$ and $\eta$. In case of backgrounds, to mitigate observed differences with collision data; respective efficiency corrections are applied CMS:2018rym ; CMS:2020uim ; CMS-PAS-JME-16-003 ; CMS:2017wtu . Table 6 lists signal and background yields after all event selection steps for the two lepton flavours, while Figures 68 show the distributions of observables related to lepton, jet, and event kinematics as indicated in Table 8.
Table 6: Event yields of signal and background processes for the two lepton flavours after applying all selection criteria for an integrated luminosity of 300 fb-1.
Final state Bkg. processes Total Bkg. Signal
$\mathrm{t\bar{t}}$ tW $\bar{\text{t}}$W $\mathrm{t\bar{t}Z}$ $\mathrm{t\bar{t}W}$ WV + jets QCD
+ jets + jets (V=~=~W or Z) multijet
ee~+ jets 18313 29544 620 825 4077 35634 89013 25
$\mu$~+ jets 69083 99413 2112 3035 13243 63309 250196 15
Reconstruction of Dynamic Radius Jets
In addition to the AK4 and AK8 jets, we reconstruct the DR anti-$\text{k}_{\text{T}}$ jets (see Section III.1) for the selected events [^4] The jet selection criteria are applied to AK4 and AK8 jets, and not to dynamic radius jets. This choice is made primarily due to the absence of optimized pile up contamination removal, jet identification working points, jet energy resolution and correction etc. for the dynamic radius jets. using the particle-flow candidate information as input. The DR algorithm is used to form two jet collections per event, corresponding to $\rm R_{0}$} values of $0.3$ and $0.7$ for narrow and fat jets, respectively. The narrow jets that do not overlap with any fat jet in the event, are considered for further analysis. The CSVv2 score of the nearest AK4 jet is assigned as the b-tagging discriminant value for each narrow jet.
Signal Event Reconstruction with Dynamic Radius Jets
It is important to understand the impact of DR clustering method on the signal event reconstruction, especially in the boosted regime. A simple study is performed for this purpose by applying the following selections on the jet substructure observables shown in Figure 5, to identify jets due to hadronically decaying boosted W or Z bosons and top quarks in the simulated signal events.
[…]
The results are presented in Table 7 in terms of the selection efficiencies of the boosted W or Z bosons and top quarks for the two clustering methods at high fat jet $\text{p}_{\text{T}}$. Higher selection efficiencies are observed for both the boosted W- or Z-like and top-like jets in the case of DR clustering method, indicating better reconstruction of the daughters of the $T$ quark in the boosted regime; thus improving the overall signal event reconstruction in this region. The adaptive approach adopted by the DR clustering (see Section III.1) helps in capturing all daughters of the boosted W or Z boson or the top quark originating from the $T$ quark decay (see Table 2) within the fat jet by optimizing the jet size.
Table 7: Selection efficiencies of boosted W- or Z-like and top-like jets for fat jet $\text{p}_{\text{T}}$ $\geqslant 900$~GeV corresponding to fixed- and dynamic radius clustering methods in simulated signal events.
Process Clustering Selection efficiency
method Boosted W or Z boson Boosted top quark
Signal Fixed-radius $38.3\%$ $46.5\%$
Dynamic radius 41.1% 47.5%
Figure 6: Distributions of lepton $\text{p}_{\text{T}}$ (upper), fat jet $\text{p}_{\text{T}}$ (middle), leading narrow jet $\text{p}_{\text{T}}$ (lower), in the $e$ + jets (left) and $\mu$ + jets (right) final states.
[…]
Figure 8: Distributions of event shape observables, namely, $10^{\text{th}}$ order Fox–Wolfram moment (upper), Sphericity (middle), and Aplanarity (lower), in the $e$ + jets (left) and $\mu$ + jets (right) final states.
V Results
V.1 Multivariate Analysis
Several observables are combined into boosted decision tree (BDT) discriminants to effectively separate signal from the collective SM backgrounds listed in Table 3. The input observables to the BDT discriminants are listed in Table 8. The most important aspects of the BDT analysis are discussed in details below.
Table 8: List and description of input observables to the BDT discriminants.
Type Observable Description
[…]
$|\eta_{\ell}|$ Lepton $|\eta|$
[…]
$|\eta_{\text{j}}|$ $|\eta|$ of the leading narrow jet that is not part of any fat jet
[…]
$|\eta_{\text{fj}}|$ Leading fat jet $|\eta|$
[…]
$\tau_{21}$ 2-to-1 subjettiness ratio of the leading fat jet
$\tau_{32}$ 3-to-2 subjettiness ratio of the leading fat jet
Event kinematics $\text{p}_{\text{T}}^{\text{miss}}$ Missing transverse momentum
[…]
$\text{H}_{\text{T}}$ Scalar $\text{p}_{\text{T}}$ sum of non-overlapping narrow and fat jets
$_{\text{jets}} N$umber of narrow jets that are not part of the fat jet
$_{\text{b-jets}} N$o. of narrow b-jets that are not part of any fat jet
[…]
S Sphericity obtained from the eigenvalues of the momentum tensor ATLAS:2012tch
A Aplanarity obtained from the eigenvalues of the momentum tensor ATLAS:2012tch
Choice of Inputs
A variety of jet substructure and event observables have been studied for this purpose, some of which are shown in Figures 48. However, a subset of these are chosen as inputs to the BDT discriminants since there are large correlations between some of the pairs. Observables that were highly correlated with each other were removed from the list of inputs to keep only the less correlated ones. For example, none of the ECFs or their ratios have been used as the inputs to the BDT discriminants because they would be highly correlated to $N$-subjettiness and their ratios. Different observables have different capabilities of discriminating the signal process from the SM backgrounds which can be broadly classified into two types, namely reducible and irreducible. The backgrounds having top quarks in the final state compose the irreducible type; while the others constitute the reducible type. Boosted decision trees are fully tree-based, human-readable structure with a finite maximum depth that selects per tree only a few from the input observables list which provide the best discrimination between signal and backgrounds. Thus, using a relatively large number of observables as inputs to the BDTs does not introduce any bias; instead it mostly increases the computational cost due to the residual correlations among the observables, even in the worst case where the performance is not significantly improved by including an observable. This strategy has been widely adopted in experimental searches such as Ref. CMS:2022fck as well as in phenomenological studies such as Ref. Ghosh:2025gue . The input observables are ranked according to their so-called “importance” to the BDT discriminants as shown in Fig. 12 in Appendix B. The “importance” of an observable is derived by counting how often that observable has been used to split decision tree nodes, and by weighting each split occurrence by the separation gain-squared it has achieved and by the number of events in the node Breiman:2017lcz .
Training Setup
The simulated events are separated into two independent sub-samples, one for training and the other for validation, in $70\%:30\%$ ratio. An event is picked randomly to belong to any one of these sub-samples. The adaptive boosting algorithm Freund:1997xna implemented within the TMVA package Hocker:2007ht has been used to train the discriminants keeping the learning rate, minimum node size, and maximum depth set to $0.5$, $2.5\%$, and $3$, respectively. Two independent trainings are performed, one for fixed-radius jets and another for DR jets, with exactly same settings. In each training, samples are selected after applying a $\text{p}_{\text{T}}$ threshold of $200$ GeV on the leading fat jet as indicated in Section IV. Each process is weighted according to its purity, defined as $\frac{N_{i}}{\sum_{i}N_{i}}$ with $i$ running over the signal and background processes, during training. Checks for overtraining are performed by comparing the respective BDT scores of signal and background between training and validation sub-samples. The corresponding Kolmogorov-Smirnov probabilities Chakravarti_Laha_Roy_1967 range between $32\%$$97\%$, indicating no overtraining. In the left column of Fig. 9, we show the BDT score distributions of signal and background corresponding to fixed-radius (upper left) and DR (lower left) algorithms for the nominal event selection.
Performance Evaluation
We evaluate the BDT scores also for events passing different thresholds on the leading fat jet $\text{p}_{\text{T}}$, and the performances of the BDT discriminants are then studied via the receiver-operator characteristic (ROC) curves for each of these thresholds separately, along with the nominal case. This exercise is primarily aimed at understanding the sensitivity in the high fat jet $\text{p}_{\text{T}}$ regime. The ROC curves corresponding to different fat jet $\text{p}_{\text{T}}$ thresholds are shown in the right column of Fig. 9 for fixed-radius (upper right) and DR (lower right) clustering methods. The BDT performance falls off less rapidly for tighter thresholds on fat jet $\text{p}_{\text{T}}$ in the case of DR jet clustering as indicated by the ROC curves in Fig. 9.
Figure 9: BDT scores (left) along with the ROC curves (right) for fixed-radius (upper) and dynamic radius (lower) jets. The BDT score distributions are obtained for the nominal fat jet $\text{p}_{\text{T}}$ threshold of $200$ GeV (see Section IV) which corresponds to the green curves on the right column on each row. The other ROC curves are obtained by applying an additional criteria on the fat jet $\text{p}_{\text{T}}$ as indicated in the legend.
V.2 Signal Sensitivity and Interpretation
The BDT discriminants are then utilized for a comparison of signal sensitivities followed by interpretations in terms of expected limits on the plane spanned by $T$ quark mass and its chromomagnetic coupling, corresponding to the two clustering approaches.
Comparative Study of Signal Sensitivities
Selection thresholds on the BDT scores are optimized based on the so-called “Punzi figure-of-merit” ($\text{FOM}_{\text{Punzi}}$) defined in Ref. Punzi:2003bu . The signal sensitivity is characterised via two different metrics, namely the signal significance defined as $\frac{S}{\sqrt{S+B}}$ and the $\text{FOM}_{\text{Punzi}}$, to check the consistency and robustness of the conclusion. The variation of these metrics are studied with the fat jet $\text{p}_{\text{T}}$ threshold for the fixed-radius and DR jet clustering methods after applying the optimized selection on the BDT scores $\geqslant 0.3$. From Figure 10, it is evident that the DR jet algorithm outperforms the fixed-radius one in the highly boosted (fat jet $\text{p}_{\text{T}}\geqslant 750$~GeV) region of the phase space. The DR algorithm shows higher sensitivity towards large $M_{T}$ values (see Fig. 3) compared to the fixed-radius algorithm, allowing for more phase space to hunt for the signal process.
Interpretation via Expected Limits
The result of the signal sensitivity study is interpreted by calculating the expected $2\sigma$ upper limits on the chromomagnetic couplings, assuming $C_{tRL}=C_{tLR}$, for the top partner masses in the range 1.4 TeV to 3 TeV. For each mass point, the expected upper limit has been obtained by varying the coupling strength such that the signal significance reaches $2\sigma$. Figure 11 shows the resulting upper limits for the three different fat jet $\text{p}_{\text{T}}$ thresholds. In these plots, the shaded regions correspond to values of the coupling excluded by our analysis at $95\%$ CL under the assumption of no signal observation. The upper limits are estimated for both fixed-radius and DR jet clustering methods. In all cases, the expected limit for the DR clustering is smaller than that for the fixed radius one, indicating that the analysis with DR jets has higher sensitivity to smaller coupling values. The improvement is more prominent for the fat jet $\text{p}_{\text{T}}$ threshold of 900 GeV, suggesting that the DR jet clustering is particularly effective for reconstructing signal events in the high fat jet $\text{p}_{\text{T}}$ regime, as indicated earlier in Section IV.
Figure 10: Comparison of $\text{FOM}_{\text{Punzi}}$ (left) and signal significance (right) between fixed-radius and dynamic radius clustering methods as a function of the fat jet $\text{p}_{\text{T}}$ threshold for the benchmark scenario defined in Eq. (7).
Figure 11: Expected $2\sigma$ upper limit on the chromomagnetic coupling $C_{tRL}=C_{tLR}$ as a function of the top partner mass. The three panels correspond to three different fat jet $\text{p}_{\text{T}}$ thresholds. For these plots, we use $\Lambda=8$~TeV, and all other parameters are kept the same as in Eq. (7). Shaded regions indicate the parameter space excluded by our analysis, assuming no signal is observed.
VI Summary and Outlook
In this work, we performed a thorough investigation of collider prospects of a vector-like top partner produced in association with a standard model top quark via a chromomagnetic coupling. The top partner is assumed to decay into the following channels: tZ, th, tg, or bW. To effectively suppress the huge QCD background, we focused on semileptonic final states, i.e., events with exactly one isolated lepton arising either from the top quark or from the heavy boson decay. Depending on the decay modes of the top partner, the final state could contain one or two fat jets along with a single lepton and b-tagged jet(s). We performed the analysis in the event category that contains a single fat jet ($\ell + 1$ fat jet + $\geqslant 1$ b-tagged jet) due to its comparatively better signal-to-background ratio.
We carried out a detailed collider analysis at the $\sqrt{s}=13$~TeV with an integrated luminosity of 300 fb-1, incorporating realistic pileup conditions. For signal events, the Madgraph5_aMC@NLO + Pythia8 + Delphes simulation chain has been employed; while for background estimation, appropriate simulated samples from CMS Open Data are utilized. To enhance signal discrimination, we utilized state-of-the-art jet substructure tools such as Soft Drop grooming, and $N$-subjettiness, along with several event shape observables like Fox-Wolfram moments, Sphericity, and Aplanarity. Multivariate discriminants based on boosted decision trees are developed for better signal-to-background separation. We then optimized a selection thresholds on these discriminants based on the Punzi figure-of-merit to improve signal sensitivity. The result is interpreted as the upper limit on the chromomagnetic coupling between the top quark and its heavy vector-like partner. Assuming the cut off scale of $\Lambda=8$ TeV, the upper limit on the coupling is found to lie within 0.2 to 1.5 for the top partner mass range of 1.4 to 3 TeV.
An important aspect of our work is a comparative performance analysis between conventional fixed-radius vs modern dynamic radius clustering of anti-$\text{k}_{\text{T}}$ jets under realistic pileup conditions. We observe that the dynamic radius approach enhances the signal sensitivity, particularly in regions with high jet transverse momentum allowing for tighter bounds on the chromomagnetic coupling. This suggests that the dynamic radius jet clustering is more effective in probing high-energy regimes compared to the conventional fixed-radius method.
There remains various scopes for further improvements. Optimization of pile up contamination removal working points as well as determining dedicated jet energy correction and resolution for the dynamic radius clustering approach can be extremely useful. This might put even tighter bounds on the couplings than the existing ones. Future directions also include exploring the use of neural network-based discriminants or adopting a hybrid strategy that combines fixed and dynamic radius clustering depending on the fat jet’s transverse momentum. Such ideas can potentially enhance the reach and robustness of the future studies.
Acknowledgments
The authors sincerely acknowledge the organizers of the WHEPP XVII, 2023 for providing a platform to meet and initiate this work. The authors acknowledge the organizers, instructors, and facilitators of the CMS Open Data Workshop at WHEPP XVII, 2023, for the necessary training to access and process CMS Open Data samples. The authors are thankful to the computing facility at IISER Kolkata for the computational and data storage resources.
Data Availability Statement
This study is based on a combination of simulated data generated using standard HEP packages and simulated data from CMS Open Data repository. The details of the simulation process that support our findings are provided within the article.
Appendix A Cross Sections and Branching Ratios for Different Mass Hypotheses
Table 9: Cross sections and branching ratios of various decay modes for different top partner masses. All other parameters are kept the same as in the benchmark point defined in Eq. (7).
$M_{T}$ [TeV] $\sigma(pp\to T\overline{t}+\overline{T}t)$ [fb] $\Gamma_{T}$ [GeV] $\mathcal{B}(T\to tZ$) $\mathcal{B}(T\to th$) $\mathcal{B}(T\to tg$) $\mathcal{B}(T\to bW$)
1.4 139.81 18.48 41.58% 9.23% 31.68% 17.52%
1.6 78.24 27.67 41.95% 9.09% 31.49% 17.47%
1.8 45.04 39.43 42.24% 9.00% 31.31% 17.45%
2.0 26.51 54.09 42.47% 8.94% 31.13% 17.45%
2.2 15.96 71.96 42.67% 8.90% 30.97% 17.46%
2.4 9.72 93.36 42.84% 8.88% 30.81% 17.47%
2.6 6.04 118.58 42.98% 8.86% 30.67% 17.49%
2.8 3.82 147.94 43.11% 8.85% 30.53% 17.51%
3.0 2.45 181.76 43.22% 8.84% 30.41% 17.53%
Appendix B Input Observables Importance for Multivariate Analysis
Figure 12: Importance of input observables to the BDT discriminants corresponding to fixed- (left) and dynamic (right) radius jet clustering.